1.    Each month, the city of Accra generates an average of 28 pounds of waste. Assume the variable is approximately normally distributed and the standard deviation is 2 pounds. If a household is selected at random, find the probability of its generating between 27 and 31 pounds of waste per month. Draw a normal distribution curve to show the required region. 


1
Expert's answer
2021-05-24T15:57:42-0400

We have given that,

"\mu= 28"

"\sigma = 2"

We have to find the probability of its generating between 27 and 31 pounds of waste per month.


"P(27<X<31)"


"= P(\dfrac{27-28}{2}<Z<\dfrac{31-28}{2})"


"= P(-0.5<Z<1.5)"


"= P(0<Z<0.5)+P(0<Z<1.5)"


"= 0.19+0.43\\=0.62"